OpenStudy (anonymous):

What is the length of the hypotenuse in a triangle who legs measure 8 inches and 6 inches?

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

8^2+6^2=c^2

OpenStudy (anonymous):

So this is the answer

OpenStudy (anonymous):

100=c^2 so solve for c and its 10

OpenStudy (anonymous):

So whats the real answer here

OpenStudy (anonymous):

10

OpenStudy (anonymous):

I just showed the steps is all

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Those steps look ruff to me

OpenStudy (anonymous):

its not that bad its one formula and you just plug in what you have and solve for the missing term

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\]

OpenStudy (anonymous):

pythagorium theorem